OpenStudy (anonymous):

find the derivative f(x)=6 3sqrt x-0.2x^10-5/x+12

OpenStudy (anonymous):

\[f(x)=6\sqrt[3]{x}-0.2x ^{10}-5/x ^{2}+12\]

OpenStudy (helder_edwin):

is this \[ \Large f(x)=6\sqrt[3]{x}-0.2x^{10}-\frac{5}{x}+12 \]

OpenStudy (helder_edwin):

do you know what is the rule for \[ \Large (x^n)'=? \]

OpenStudy (anonymous):

No haven't learned it yet

OpenStudy (helder_edwin):

OK the rule says \[ \Large (x^n)'=nx^{n-1} \]

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

now to find the derivative you have to work on every summand separatedly. do you know this?

OpenStudy (anonymous):

No :(

OpenStudy (helder_edwin):

the most basic rules of derivatives are: \[ \Large (k)'=0\qquad\text{$k$ is a constant} \] \[ \Large (u+v)'=u'+v' \] \[ \Large (uv)'=u'v+uv' \] \[ \Large \left(\frac{u}{v}\right)'=\frac{u'v-uv'}{v^2} \]

OpenStudy (helder_edwin):

and \[ \Large (cu)'=c\cdot u'\qquad\text{$c$ is a constant} \]